A company took to linear regression analysis to estimate their sales for the coming year,based on past sales, the company obtained the following equation in the format Y=a+bx: Y=0,5x+20 units in millions. If the equation was for the 10th period in the series, what would the estimate be for the coming year period 11?
1
Expert's answer
2020-08-12T18:20:57-0400

y=0,5x + 20 is for the period of 10th.

Therefore;

y=0,5(10) +20

y=R25.0 million

For the period of 11th,

y=0,5x + 20

y= 0,5(11) +20

y=5,5 + 20

y=225,5

R 25,5.0 million